# Closed-Form Power and Sample Size Calculations for Bayes Factors

### v1

@@ 6.2 Normal moment priors
Nonlocal priors are prior distributions that have no probability density/mass at the null value. They were introduced to allow evidence for the null to accumulate more quickly if it is indeed true (Johnson and Rossell,, 2010). A convenient class of nonlocal priors are normal moment priors, which have a density of the form $\mathrm{NM}(\theta\mid\theta_{0},\tau)=\mathrm{N}(\theta\mid\theta_{0},\tau^{2})\times(\theta-\theta_{0})^{2}/\tau^{2}$ with location $\theta_{0}$ and spread $\tau$. Figure 5 shows three examples of normal moment priors. We see that the distributions have a density of zero at the null value $\theta_{0}=0$, encoding the assumption that the null is the least likely parameter value under the alternative. Normal moment priors have two modes at ±$\pm\tau\sqrt{2}$, which gives a convenient way to elicit a prior, as $\tau$ may, for example, be set so that the mode equals two parameter values deemed most plausible under the alternative.
@@
We will now extend power and sample size calculation to Bayes factors with nonlocal moment analysis priors and normal design priors. The Bayes factor based on a normally distributed parameter estimate $\hat{\theta}\mid\theta\sim\mathrm{N}(\theta,\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n)$ that contrasts $H_{0}\colon\theta=\theta_{0}$ to $H_{1}\colon\theta\neq\theta_{0}$ with prior $\theta\mid H_{1}\sim\mathrm{NM}(\theta_{0},\tau)$ assigned to $\theta$ under $H_{1}$ is given by
@@
with $\mathrm{W}_{0}$ the principal branch of the Lambert W function, see Appendix E for details. Based on the power function (14), numerical root-finding can again be used to determine the sample size such that a specified target power is achieved. Methods for power and sample size calculations for normal moment prior Bayes factors are both implemented in our R package (functions nmbf01 and powernmbf01).
Figure 6: Power and sample size calculations for standardized mean difference example from Schönbrodt and Wagenmakers, (2018) based on normal moment prior Bayes factor. The normal moment analysis prior has modes at a medium effect size ($\tau=0.5/\sqrt{2}$). The power is computed under either a point prior at a medium effect size (green), a normal prior that incorporates additional parameter uncertainty (blue), or under the null hypothesis (orange).
Figure 6 shows again power and sample size calculations for the example from Schönbrodt and Wagenmakers, (2018) considered earlier. This time, however, we specify a normal moment analysis prior with modes at a medium standardized mean difference effect size of $\theta=\pm 0.5$, which translates to a prior spread parameter of $\tau=0.5/\sqrt{2}$. Using the same design priors as before, we can see that the sample sizes required to achieve a target power of 95% are larger than when using the normal analysis prior as in Section 5.2. For example, under the point design prior a sample size of $n=302$ per group is required while only $n=153$ were required with the normal analysis prior (see Figure 4). At the same time, the sample size required to find evidence in favor of the null hypothesis with a target power of 95% (bottom plot) is drastically reduced ($n=997$) compared to the normal analysis prior ($n=6691$). This is expected since the evidence for a true null hypothesis accumulates faster with normal moment priors (Johnson and Rossell,, 2010).
@@ 7 Discussion
A clear limitation of our methodology is the asymptotic normality assumption. This assumption may be inappropriate for certain data or parameter types, and may lead to an underappreciation of uncertainty and consequently an underestimation of sample size. Simulation-based methods do not have this shortcoming, as they can be tailored to any data distribution and analysis method. Nevertheless, simulation methods may be intimidating or too advanced for research workers, in which case we believe it is better to do an approximate calculation than no calculation at all. One avenue for future work might be to extend closed-form power and sample size calculations to more specific settings not considered here, such as, binary outcomes. Another limitation is the types of Bayes factors that we considered for the analysis, which is limited to univariate parameters with normal, normal moment, $t$, or point priors under the alternative. Our work could be extended to ANOVA or regression settings with multivariate parameters and/or to other prior distributions. We also did not consider ‘open-ended’ sequential designs, where data are collected continuously until compelling evidence for one of the competing hypotheses is found (Wald,, 1947). The sequential approach is particularly attractive for Bayes factor inference as design and analysis prior distributions can be updated based on the accumulating data. Researchers can then make informed decisions about whether or not it is worthwhile to continue collecting data or to stop (Stefan et al.,, 2024). For these purposes, it would be interesting to consider the Bayes factor indexed by the sample size as a stochastic process, and study its properties. Finally, in situations where researchers have only a fixed sample size at their disposal, it may be interesting to use a ‘reverse-Bayes’ approach (Held et al.,, 2022). That is, one may ‘go backwards’ and determine the design prior required to achieve a desired target power. Researchers can then reason whether or not this ‘reverse Bayes design prior’ is scientifically plausible, and whether or not they should undertake the study based on their limited resources.
@@ Acknowledgments
We thank Angelika Stefan and František Bartoš for valuable comments on drafts of the manuscript. The acknowledgment of these individuals does not imply their endorsement of the paper.
@@
Code and data to reproduce our analyses are openly available at https://github.com/SamCH93/bfpwr. A snapshot of the repository at the time of writing is available at https://doi.org/10.5281/zenodo.12582277. We used the statistical programming language R version 4.4.0 (2024-04-24) for analyses (R Core Team,, 2024) along with the BFDA (Schönbrodt and Stefan,, 2019), lamW (Adler,, 2015), xtable (Dahl et al.,, 2019), and knitr (Xie,, 2024) packages.
@@ Appendix A The bfpwr R package
Our R package can be installed from GitHub with the R command remotes::install_github(repo = "SamCH93/bfpwr", subdir = "package") (requires the remotes package). We plan to submit the package to CRAN in the near future. The workhorse function of our R package is powerbf01. It is inspired by the power.t.test function from the stats package, with which many user will be familiar. As power.t.test, the function powerbf01 assumes that the data are continuous and that the parameter of interest is either a mean or a (standardized) mean difference. The functions pbf01 and nbf01 are more general and can be used for any approximately normally distributed parameter estimate with approximate variance $\mathrm{Var}(\hat{\theta})=\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n$, although users have to specify the unit variance $\sigma^{2}_{\scriptscriptstyle\hat{\theta}}$ themselves. Similar functions exist also for $t$-test Bayes factors (powertbf01) and normal moment prior Bayes factors (powernmbf01). The following code chunk illustrate how powerbf01 can be used.
@@
k <- 1/6 # BF threshold
null <- 0 # null value
sd <- 1 # standard deviation of one observation
pm <- null # analysis prior mean set to the null value
psd <- sqrt(2) # analysis prior sd set to sqrt(2)
type <- "two.sample" # two-sample test
@@
dpm <- 0.5 # design prior mean equal to large SMD effect size
dpsd <- 0.1 # design prior sd to incorporate parameter uncertainty
@@
##       n is number of *samples per group*
##       sd is standard deviation of one sample (assumed equal in both groups)
@@ Appendix D Sample size for Bayes factor with local normal prior
from which we get the sample size
@@
Appendix E Power with normal moment prior
@@
Since the argument of the Lambert W function is real and always non-negative, only the principal branch $\mathrm{W}_{0}$ can satisfy the inequality (Corless et al.,, 1996). Assuming a $\hat{\theta}\mid n,\mu_{d},\tau_{d}\sim\mathrm{N}(\mu_{d},\tau^{2}_{d}+\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n)$ distribution induced by a normal design prior, we can rearrange the inequality (17) and obtain the power function (14).

### v2

@@ 6.2 Normal moment priors
Non-local priors are prior distributions that have zero probability density/mass at the null value. They were introduced to allow evidence for the null to accumulate more quickly if it is indeed true (Johnson and Rossell, , 2010). A convenient class of non-local priors are normal moment priors, which have a density of the form $\mathrm{NM}(\theta\mid\theta_{0},\tau)=\mathrm{N}(\theta\mid\theta_{0},\tau^{2})\times(\theta-\theta_{0})^{2}/\tau^{2}$ with location $\theta_{0}$ and spread $\tau$. Figure 5 shows three examples of normal moment priors. We see that the distributions have a density of zero at the null value $\theta_{0}=0$, encoding the assumption that the null is the least likely parameter value under the alternative. Normal moment priors have two modes at ±$\pm\tau\sqrt{2}$, which gives a convenient way to elicit a prior, as $\tau$ may, for example, be set so that the modes equals two parameter values deemed most plausible under the alternative.
@@
We will now extend power and sample size calculation to Bayes factors with non-local moment analysis priors and normal design priors. The Bayes factor based on a normally distributed parameter estimate $\hat{\theta}\mid\theta\sim\mathrm{N}(\theta,\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n)$ that contrasts $H_{0}\colon\theta=\theta_{0}$ to $H_{1}\colon\theta\neq\theta_{0}$ with prior $\theta\mid H_{1}\sim\mathrm{NM}(\theta_{0},\tau)$ assigned to $\theta$ under $H_{1}$ is given by
@@
with $\mathrm{W}_{0}$ the principal branch of the Lambert W function, see Appendix F for details. Based on the power function (15), numerical root-finding can again be used to determine the sample size such that a specified target power is achieved. Methods for power and sample size calculations for normal moment prior Bayes factors are implemented in our R package (functions nmbf01 and powernmbf01).
Figure 6: Power and sample size calculations for standardized mean difference example from Schönbrodt and Wagenmakers, (2018, p. 133) based on normal moment prior Bayes factor. The normal moment analysis prior has modes at a medium effect size ($\tau=0.5/\sqrt{2}$). The power is computed under either a point prior at a medium effect size (green), a normal prior that incorporates additional parameter uncertainty (blue), or under the null hypothesis (orange).
Figure 6 shows again power and sample size calculations for the example from Schönbrodt and Wagenmakers, (2018) considered earlier. This time, however, we specify a normal moment analysis prior with modes at a medium standardized mean difference effect size of $\theta=\pm 0.5$, which translates to a prior spread parameter of $\tau=0.5/\sqrt{2}=0.35$. Using the same design priors as before, we can see that the sample sizes required to achieve a target power of 95% are larger than when using the normal analysis prior as in Section 5.2. For example, under the point design prior a sample size of $n=302$ per group is required while only $n=153$ were required with the normal analysis prior (see Figure 4). At the same time, the sample size required to find evidence in favor of the null hypothesis with a target power of 95% (bottom plot) is drastically reduced ($n=997$) compared to the normal analysis prior ($n=6691$). This is expected since the evidence for a true null hypothesis accumulates faster with normal moment priors (Johnson and Rossell, , 2010).
@@ 7 Discussion
An important issue in any Bayesian design and analysis is the choice of prior distributions. While there are differences between point and normal priors in terms of the closed-form availability of sample sizes (see Table 1), we believe that the choice between them should be guided only by prior knowledge and the questions researchers wish to answer. Point analysis priors seem more appropriate when researchers have precise hypotheses to test, they can also be motivated from a likelihood perspective, and they are easier to communicate to a non-statistical audience. In contrast, normal analysis priors can incorporate uncertainty and may therefore be more appropriate for exploratory settings where little is known about the phenomenon under investigation (Held and Ott, , 2018). Similar considerations apply to the choice of the design prior. For example, a point design prior may be specified at a minimally relevant parameter value if such a value can be formulated. On the other hand, if such a value is difficult to formulate but data from a previous study are available, a normal design prior that takes the previous estimate and its standard error as the mean and standard deviation may be a reasonable way to incorporate prior knowledge and uncertainty.
@@
A clear limitation of our methodology is the asymptotic normality assumption. This assumption may be inappropriate for certain data or parameter types, and may lead to an underappreciation of uncertainty and consequently an underestimation of sample size. Simulation-based methods do not have this shortcoming, as they can be tailored to any data distribution and analysis method. Nevertheless, simulation methods may be intimidating or too advanced for research workers, in which case we believe it is better to do an approximate calculation than no calculation at all. One avenue for future work might be to extend closed-form power and sample size calculations to more specific settings not considered here, such as, binary outcomes.
Another limitation is the types of Bayes factors that we considered for the analysis, which is limited to univariate parameters with normal, normal moment, $t$, or point priors under the alternative. However, after publishing an initial version of this article on the arXiv preprint server, Wong and Tendeiro, (2024) have built on our results and extended the root-finding approach from Section 6 to more flexible design prior distributions and $t$ likelihoods. Nevertheless, the approach could even be further extended to ANOVA or regression settings.
Furthermore, in the case of two-group parameters, we have assumed equal allocation of observations between the two groups, but unequal allocation may sometimes be desirable. For example, it might be easier to recruit patients for a study if they have a higher chance of receiving a new treatment than placebo, as participation will then be more attractive. In the case of (standardized) mean difference parameters, this is usually addressed by introducing an allocation ratio $r=n_{2}/n_{1}$ and changing the sample size per group $n$ based on equal allocation to $n_{1}=(2n)/(1+r)$ and $n_{2}=(2n)/(1+1/r)$, see e.g., Kieser, (2020, p. 21). While this adjustment has been developed for frequentist sample size determination, it is equally applicable to the Bayes factor methods considered here. However, for other parameter types, such as log hazard ratios, this is more complicated (Kieser, , 2020, p.74-75).
We also did not consider sequential designs, where data are collected continuously until compelling evidence for one of the competing hypotheses is found (Wald, , 1947), or the sample size is ‘recalculated’ after an interim analysis (see part III in  Kieser, , 2020). The sequential approach is particularly attractive for Bayes factor inference as design and analysis prior distributions can be updated based on the accumulating data. Researchers can then make informed decisions about whether or not it is worthwhile to continue collecting data or to stop (Stefan et al., , 2024). For these purposes, it would be interesting to consider the Bayes factor indexed by the sample size as a stochastic process, and study its properties.
Finally, in situations where researchers have only a fixed sample size at their disposal, it may be interesting to use a ‘reverse-Bayes’ approach (Held et al., , 2022) and determine the design prior required to achieve a desired target power. Researchers can then reason whether or not this prior is scientifically plausible and they should undertake the study based on their limited resources. This is similar to classical power analysis, where one can determine the value of the true parameter needed to achieve a given power for a given sample size, and then reason about whether or not this true parameter is realistic. In both cases, however, it is important that the focus is on reasoning about the plausibility of the design prior/parameter, and researchers should be careful not to give the impression that this choice has been made a priori.
@@ Acknowledgments
We thank Angelika Stefan and František Bartoš for valuable comments on drafts of the manuscript. We thank Tsz Keung Wong for helpful suggestions related to our R package. We thank two anonymous reviewers and an associate editor for constructive comments. The acknowledgment of these individuals does not imply their endorsement of the paper.
@@
Code and data to reproduce our analyses are openly available at https://github.com/SamCH93/bfpwr. A snapshot of the repository at the time of writing is available at https://doi.org/10.5281/zenodo.12582277. We used the statistical programming language R version 4.4.1 (2024-06-14) for analyses (R Core Team, , 2024) along with the BFDA (Schönbrodt and Stefan, , 2019), lamW (Adler, , 2015), xtable (Dahl et al., , 2019), and knitr (Xie, , 2024) packages.
@@ Appendix A The bfpwr R package
Our R package can be installed by running install.packages("bfpwr") in an R session, the development version can be found on GitHub (https://github.com/SamCH93/bfpwr). The workhorse function of our R package is powerbf01. It is inspired by the power.t.test function from the stats package, with which many user will be familiar. As power.t.test, the function powerbf01 assumes that the data are continuous and that the parameter of interest is either a mean or a (standardized) mean difference. The functions pbf01 and nbf01 are more general and can be used for any approximately normally distributed parameter estimate with approximate variance $\mathrm{Var}(\hat{\theta})=\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n$, although users have to specify the unit standard deviation $\sigma_{\scriptscriptstyle\hat{\theta}}$ themselves. Similar functions exist also for $t$-test Bayes factors (powertbf01) and normal moment prior Bayes factors (powernmbf01). The following code chunk illustrate how powerbf01 can be used.
@@
k <- 1/6 # set BF˙01 threshold to 1/6
null <- 0 # set null value to 0
sd <- 1 # set standard deviation of one observation to 1
pm <- 0 # set analysis prior mean to 0
psd <- sqrt(2) # set analysis prior SD set to sqrt(2)
type <- "two.sample" # set test to two-sample
@@
dpm <- 0.5 # set design prior mean equal to medium SMD effect size
dpsd <- 0.1 # set positive design prior SD to incorporate parameter uncertainty
@@
##       n is number of *observations per group*
##       sd is standard deviation of one observation (assumed equal in both groups)
@@ Appendix D Sample size for Bayes factor with local normal prior
from which we obtain the sample size
@@
Appendix E Simulation-based evaluation of the bfpwr package
Figure 7 shows a simulation-based evaluation of the power and sample size calculation methods for the Bayesian $z$-test as implemented in our bfpwr R package. The values for the design and analysis prior means were chosen to represent conventions for no (0), small (0.2), medium (0.5), and large (0.8) standardized mean differences (Cohen, , 1992). The null and alternative hypotheses were defined as $H_{0}\colon\theta=0$ against $H_{1}\colon\theta\neq 0$. The standard deviations were chosen to include point and normal priors. For each combination of analysis/design prior mean/standard deviation, the sample size to obtain a Bayes factor equal or below $k=1/10$ with a target power of 80% was computed (shown at the top of each plot). This sample size along with the design prior was subsequently used to simulate 50’000 standardized mean difference parameter estimates based on which 50’000 Bayes factors were computed. The power was then estimated from the proportion of Bayes factors equal or below the level $k=1/10$. Note that for certain design/analysis prior combinations, it is impossible to achieve the target power with a finite sample size. In this case an “x” is shown in the plot.
Figure 7: Simulation-based evaluation of power and sample size calculations related to the Bayesian $z$-test as implemented in the bfpwr package. The top of each plot shows the sample size to obtain a Bayes factor equal or below $k=1/10$ with a target power of 80% for the corresponding combination of analysis and design prior (an “x” is shown if the target power is impossible to achieve for a given condition). 50’000 Bayes factors were then simulated based on this sample size, from which then the power was empirically estimated.
We can see that in all conditions, the simulation-based estimate of the power closely spreads around the target power of 80%. The maximally observed discrepancy is 0.56% while the median discrepancy is 0.14%. This suggests that the power and sample size calculation methods work as intended.
Appendix F Power with normal moment prior
@@
Since the argument of the Lambert W function is real and always non-negative, only the principal branch $\mathrm{W}_{0}$ can satisfy the inequality. Assuming a $\hat{\theta}\mid n,\mu_{d},\tau_{d}\sim\mathrm{N}(\mu_{d},\tau^{2}_{d}+\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n)$ distribution induced by a normal design prior, we can rearrange the inequality (18) and obtain the power function (15).